OpenStudy (anonymous):

Write an equation of the line that passes through point p and is parallel to the line with the given equation. Problem: P (4,0), -x + 2y=12

OpenStudy (anonymous):

First get the equation you were given into slope intercept form. In other words, solve for y

OpenStudy (anonymous):

It would be y= 1/2 + 6

OpenStudy (anonymous):

I dont know what to do after that

OpenStudy (anonymous):

Alright, parallel lines have the same slope. So these two lines have the same slope...do you know what number in that equation would be the slope?

OpenStudy (anonymous):

It would be 1/2

OpenStudy (anonymous):

Yep. and you have a point the line goes through so you can do point slope form which is: y-y1=m(x-x1) ever seen that?

OpenStudy (anonymous):

I dont understand

OpenStudy (anonymous):

Thats the formula for point slope form which when simplified will give you the equation you want. (x1, y1) is the point you were given and m is the slope, so just plug those in

OpenStudy (anonymous):

y-0=1/2(x-4)

OpenStudy (anonymous):

I get it now!!! Thank you so much!!

OpenStudy (anonymous):

Did you simplify the equation?

OpenStudy (anonymous):

When you simplify you get y=1/2x-2 and that is your final answer

OpenStudy (anonymous):

Yep thanks

OpenStudy (anonymous):

You're welcome.